OpenStudy (anonymous):

F(X)= 5X^2+6 G(X)=X-4 FIND (gof)(x) algebra 2 need help please!

OpenStudy (anonymous):

This is just composition of functions. Basically if you have two functions, which ever one is second gets replace by all the x's Look here for a better explanation though. http://www.mathsisfun.com/sets/functions-composition.html

OpenStudy (anonymous):

.. yea I need help of going through the problem..

OpenStudy (anonymous):

*This is another way of writing it however* \[\Large fog=(f(g(x))\] Basically if there was f=x^2 and g=3x+5 then \[\Large f(g(x))=(3x+5)^2\] because the x is being square and in this case g(x) is what we call x and then if we did it the other way around \[\Large g(f(x))=3(x^2)+5=3x^2+5\]

OpenStudy (anonymous):

Is the answer 25x^2 -100x+306

OpenStudy (anonymous):

But lets see in this example, every where we see an x in the F function, we're gonna replace it with whatever g(x) is \[\Large F(G(x))= 5(x-4)^2+6=5(x^2-8x+16)+6=\]

OpenStudy (anonymous):

\[\Large 5x^2-40x+80+6~=~~~~~5x^2-40x+86\]

OpenStudy (anonymous):

ooh..

OpenStudy (anonymous):

Yep, just message me if you need anymore help!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

YW